OpenStudy (anonymous):

h(t)=(t+1)^2/3(2t-1)^3 h'(t)=(t+1)^2/3*3(2t^2-1)^2*4t+(2t^2-1)^3*2/3(t+1)^-1/3 I am stuck right here please help! Using the chain rule I have done ok. It's the Algebra that is killing me!

jimthompson5910 (jim_thompson5910):

There's not much else you can do here. I guess you could try to simplify, but I don't think any amount of algebra will actually simplify this. It will just rewrite it in another form really.

OpenStudy (anonymous):

You can factor

jimthompson5910 (jim_thompson5910):

yeah I guess you could factor out a t+1 term

jimthompson5910 (jim_thompson5910):

if you factor out say (t+1)^(2/3), then you divide (t+1)^2/3*3(2t^2-1)^2*4t by (t+1)^(2/3) and you also divide (2t^2-1)^3*2/3(t+1)^-1/3 by (t+1)^(2/3)

OpenStudy (anonymous):

But see you would have to factor 2/3(t+1)^1/3 and (2t^2-1)^2

OpenStudy (anonymous):

I'm hoping Perl is got a nice solid answer...

OpenStudy (perl):

one moment, i might try something different

OpenStudy (anonymous):

I see what Jim was sort of saying. What program do you use?

OpenStudy (perl):

i noticed in your question, a square appears in the derivative. is h(t)=(t+1)^2/3(2t-1)^3 or h(t)=(t+1)^2/3(2t^ `2`-1)^3

OpenStudy (anonymous):

using the chain rule it is converted to (4t)

jimthompson5910 (jim_thompson5910):

oh that's a good point @perl, I didn't notice that til just now

OpenStudy (anonymous):

So what does the square do?

OpenStudy (anonymous):

I'm sorry I am missing that point

OpenStudy (anonymous):

h(t)=(t+1?13(2t2 - 1)3 => h'(t)= (t+1?/3 . 3(2t2 - 1)2 . 4t+(2t2 -1)3 . +1)-113 = +1)-113(2t2 - 1)2[18t(t+1)+(2t2- 1)] =i(t+1)-113(2t2- 1)2(20t2+18t-1)

OpenStudy (anonymous):

I'm sorry that came out ugly

OpenStudy (perl):

$$ \large {h(t) = (t+1)^\frac{2}{3}(2t\color{red}{^2}-1)^3 \\ \large { h'(t)=(t+1)^{2/3}\cdot 3(2t^2-1)^2~4t+\frac{2}{3}(t+1)^{-1/3} (2t^2-1)^3\cdot \\ =\frac{2}{3}(2t^2-1)^2(t+1)^{-\frac{1}{3}}[~(t+1)^{\frac{2}{3}+\frac{1}{3} }\cdot 2\cdot 9 t +(2t^2 -1)] }} $$

OpenStudy (perl):

any questions so far?

OpenStudy (anonymous):

I'm following you!

OpenStudy (anonymous):

Well why are you adding 2/3+1/3?

OpenStudy (perl):

try distributing

OpenStudy (perl):

if we go backwards and distribute we have: \(\color{blue}{\text{Originally Posted by}}\) \[ { \frac{2}{3}(2t^2-1)^2(t+1)^{-\frac{1}{3}}[~(t+1)^{\frac{2}{3}+\frac{1}{3} }\cdot 2\cdot 9 t +(2t^2 -1)]\\\text{ }\\ =\\\text { }\\ \frac{2}{3}(2t^2-1)^2(t+1)^{-\frac{1}{3}}\cdot ~(t+1)^{\frac{2}{3}+\frac{1}{3} }\cdot 2\cdot 9 t +\frac{2}{3}(2t^2-1)^2(t+1)^{-\frac{1}{3}}(2t^2 -1)] \\\text { }\\=\\\text { }\\ \frac{2}{3}\cdot 2\cdot 9 t~(2t^2-1)^2 ~(t+1)^{-\frac{1}{3} +\frac{2}{3}+\frac{1}{3} } +\frac{2}{3}(2t^2-1)^{2+1}(t+1)^{-\frac{1}{3}} \\\text { }\\=\\\text { }\\ 3\cdot 4t~(2t^2-1)^2 ~(t+1)^{\frac{2}{3} } +\frac{2}{3}(2t^2-1)^{3}(t+1)^{-\frac{1}{3}} \\\text { }\\\text{so they are equivalent }\\\text { }\\ } \] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (nincompoop):

you can only use the chain rule on the denominator